OpenStudy (newwar):

What are the x- and y- intercepts for the line y+8=4(x−5)? Enter your answers as a whole number in the boxes. The x-intercept is The y-intercept is

OpenStudy (mathmale):

What have you done in the past to determine x- and y-intercepts? Given y+8=4(x−5), to find the x-intercept, let y=0. Solve for x. Write the result in the form (x,0). Please show your work.

OpenStudy (newwar):

i think the x intercept is 7 and the y is 28

OpenStudy (newwar):

-28 sorry

OpenStudy (taylor1048):

what you need to do is substitute the numbers for x and y and see if they equal the same because the equation is y+8=4(x-5)

OpenStudy (triciaal):

do you need more help or did you forget to close?

OpenStudy (newwar):

yes i do @triciaal

OpenStudy (triciaal):

the x-intercept is the value of x when Y=0

OpenStudy (triciaal):